# Reasoning Language Models: A Blueprint

### v3

@@ D.3.1 Process-Based Q-Value Modeling
For q-value models the same discussions about reward modeling apply here since the models are trained very similar. This is why omit it here.
@@ D.3.2 The Difference between Value and Q-Value Models
The difference of VMs and QVMs can be easily shown in how they are used in the evaluation processes of an MCTS algorithm. QVMs predict $\hat{Q}_{\varphi}(s_{t},a_{t})$, which evaluates the action $a_{t}$ taken in state $s_{t}$ that deterministically transitions to $s_{t+1}$. Thus, the value $\hat{Q}(s_{t},a_{t})$ is used to evaluate adding the node $s_{t+1}$ to the tree. On the other hand, for VMs, adding a node $s_{t+1}$ to the tree is determined by $\hat{V}(s_{t+1})=\frac{1}{\gamma}\hat{Q}_{\varphi}(s_{t},a_{t})$, where $\gamma$ is the discount factor.
This distinction is making the training processes different. Note that $s_{t}\frown a_{t}=s_{t+1}$. For QVMs, the training tuples are $((s_{t},a_{t}),\hat{Q}(s_{t},a_{t}))=(s_{t+1},\hat{Q}(s_{t},a_{t}))$ due to the deterministic transition. For VMs, the corresponding training tuples are $(s_{t+1},\hat{V}(s_{t+1}))$. Since we propose training VMs on terminal rewards for terminal states instead of assigning a label of 0, VMs and QVMs become equivalent under the following transformation for any $t\in\{0,\ldots,T\}$ for evaluating adding node $s_{t+1}$:

### v4

@@ D.3.1 Process-Based Q-Value Modeling
The considerations about reward modeling apply to q-value models as well since these models are trained very similar, so we omit their discussion here.
@@ D.3.2 The Difference Between Value and Q-Value Models
The difference of VMs and Q-VMs can be easily shown in how they are used in the evaluation processes of an MCTS algorithm. Q-VMs predict $\hat{Q}_{\varphi}(s_{t},a_{t})$, which evaluates the action $a_{t}$ taken in state $s_{t}$ that deterministically transitions to $s_{t+1}$. Thus, the value $\hat{Q}(s_{t},a_{t})$ is used to evaluate adding the node $s_{t+1}$ to the tree. On the other hand, for VMs, adding a node $s_{t+1}$ to the tree is determined by $\hat{V}(s_{t+1})=\frac{1}{\gamma}\hat{Q}_{\varphi}(s_{t},a_{t})$, where $\gamma$ is the discount factor.
This distinction is making the training processes different. Note that $s_{t}\frown a_{t}=s_{t+1}$. For Q-VMs, the training tuples are $((s_{t},a_{t}),\hat{Q}(s_{t},a_{t}))=(s_{t+1},\hat{Q}(s_{t},a_{t}))$ due to the deterministic transition. For VMs, the corresponding training tuples are $(s_{t+1},\hat{V}(s_{t+1}))$. Since we propose training VMs on terminal rewards for terminal states instead of assigning a label of 0, VMs and Q-VMs become equivalent under the following transformation for any $t\in\{0,\ldots,T\}$ for evaluating adding node $s_{t+1}$: